OpenStudy (candycane):

Ive read this problem several times and I dont quiet understand. You have 324 ft^2 of fabric to make a circular play parachute for kids. What is the radius of the largest parachute you can make? Round to the nearest tenth of a foot.

OpenStudy (anonymous):

\[ \pi r^2 \leq 324\\ r^2 \leq \frac {324}\pi\\ r \leq \sqrt{ \frac {324}\pi }=\frac {18}{\sqrt\pi} \]

OpenStudy (anonymous):

\[ \frac{18}{\sqrt{\pi }}=10.1554 \]

OpenStudy (candycane):

Im not sure I understand

OpenStudy (candycane):

Can some one please elaborate?

OpenStudy (anonymous):

A circular play parachute just looks like a huge circle when it's on the ground. The area of a circle is (pi)(radius)^2. They're telling you the max area is 324 square feet. pi (pronounced pie) is a number that equals 3.14159265.... It's a button on your calculator. So now you have an equation with only one unknown, the radius. Now look at elissaab's calculations. The biggest parachute that you could make would have a radius of 10.155 feet.

OpenStudy (anonymous):

It says to round to the nearest 10th of a foot, so your answer would be 10.2 feet.